OpenStudy (anonymous):

2. For the system below: (a) Solve the system by the linear combination method. Show all work carefully. (b) Check your answer. a-2b=2 a+b=2(b+3)

OpenStudy (wolf1728):

a-2b=2 a+b=2(b+3) I'd say the first step would be to get the second equation into proper form. (Do the multiplication).

OpenStudy (wolf1728):

Can you do the multiplication?